# A coisotropic embedding theorem for pre-multisymplectic manifolds

### v1

@@ 3.2 Proof
This paper has established a coisotropic embedding theorem for pre-multisymplectic manifolds. We constructed a suitable multisymplectic thickening of any given pre-multisymplectic manifold such that the original manifold can be embedded as a coisotropic submanifold. This embedding facilitates a more accessible finite-dimensional approach to regularization techniques, which traditionally require handling infinite-dimensional spaces.
Building on these results, several directions for future research emerge.
@@
We will apply the coisotropic embedding theorem proved to Field Theories, particularly in constructing a Poisson bracket on the space of solutions, as we already did in [CDI+24, CDI+22b, CDI+22c]. We will study the relation between the two approaches and the potential advantages of the one based on the multisymplectic coisotropic embedding theorem proved in this manuscript.
@@
Implicit partial differential equations can often be formulated using pre-multisymplectic structures. The regularization method introduced in this paper allows for the definition of a multisymplectic structure on an extended manifold, enabling (under some conditions that have to be studied) the formulation of an explicit PDE. Future work will investigate:
  • The correspondence between solutions of the original implicit PDE and the explicit PDE on the extended manifold.
  • The conditions under which the integrability of the explicit PDE ensures the integrability of the original system.

### v2

@@ 3.1 Equations of motion
In this paper, we proved a coisotropic embedding theorem for pre-multisymplectic manifolds. We constructed a suitable multisymplectic thickening of any given pre-$k$-plectic manifold in which it is embedded as a (k1)(k-1)-coisotropic submanifold. It is the bundle of transversal (k1)(k-1)-forms on the pre-multisymplectic manifold associated with some connection. This embedding may be employed for a finite-dimensional approach to the regularization techniques listed in the introduction, which traditionally require handling infinite-dimensional spaces.
Building on these results, several directions for future research emerge, as we are going to explain.
@@
We will apply the coisotropic embedding theorem proved to Field Theories, particularly in constructing a Poisson bracket on the space of solutions, as we already did in [CDI+24, CDI+22b, CDI+23]. We will study the relation between the two approaches and the potential advantages of the one based on the multisymplectic coisotropic embedding theorem proved in this manuscript.
@@
Implicit partial differential equations can often be formulated using pre-multisymplectic structures. The regularization method introduced in this paper allows for the definition of a multisymplectic structure on an extended manifold, enabling (under some conditions that must be studied [^2] A first, obvious, one is: if the pre-multisymplectic manifold one starts with is a first-order jet bundle (or its reduced dual), under which conditions the multisymplectic thickening is a first-order jet bundle as well?) the formulation of an explicit PDE. Future work will investigate:
  • The correspondence between solutions of the original implicit PDE and the explicit PDE on the multisymplectic thickening.
  • The conditions under which the integrability of the explicit PDE ensures the integrability of the implicit one.
@@
Acknowledgements
I am indebted to prof. A. Ibort for making me aware of the existence of the coisotropic embedding theorem and its potential applications as well as for the huge amount of inspiring discussions of the last years. I am also indebted to prof. G. Marmo for encouraging me to think of alternative ways to depict the geometry behind Gotay’s coisotropic embedding theorem. I acknowledge financial support from Next Generation EU through the project 2022XZSAFN – PRIN2022 CUP: E53D23005970006. I am a member of the GNSGA (Indam).
Conflict of interests/Competing interests
The author has no conflict of interests or competing interests to disclose.